Solve for x and y in the following problems:
a. x +y=10, 3x + y =12
b. 2x + 5y = 19, 3x + 3y = 15
c. 4x + y = 22, 2x + 3y = 16
d. 12x + y = 174, 8x -2y = 36
Show work so I can understand how you would find the final answer. Thank You!

Answer by Alan3354(69443)

Solve for x and y in the following problems:
a. x +y=10, 3x + y =12
b. 2x + 5y = 19, 3x + 3y = 15
c. 4x + y = 22, 2x + 3y = 16
d. 12x + y = 174, 8x -2y = 36
------------------------------
There's more than one way to do these.
a. In this system, the coefficient of y is the same. That makes elimination easy.
x + y=10
3x +y =12
------------ Subtract
-2x = -2
x = 1
--------
Then sub into either eqn to find y.
y = 9
------------
b. Divide the 2nd eqn by 3, then multiply it by 2. Then eliminate x as in part a.
2x + 5y = 19
3x + 3y = 15 --> x + y = 5 --> 2x + 2y = 10
--------------
2x + 5y = 19
2x + 2y = 10
------------- Subtract, etc.
c. Multiply the 2nd eqn by 2, then use the same method.
4x + y = 22
2x + 3y = 16
---------------
d. Divide the 2nd eqn by 2, then add them, etc.
12x + y = 174
8x - 2y = 36
--------------
email me via the thank you note for assistance, or to check your work.